OpenStudy (anonymous):

Can someone please tell me how I would go about finding the pH of 0.001 M solution of HCl

OpenStudy (jfraser):

HCl is a strong acid, so all of teh molecules will dissociate into H+ ions (and Cl- ions that don't matter here). pH is:\[pH = -\log[H^{+1}] = -\log(1*10^{-3}) = 3\]

OpenStudy (anonymous):

Thank you!